a bag contains 4 balls. two balls are drawn at random and found to be white. what is the probability that that that the bag contained all 4 white balls ?

Grade:

1 Answers

shael shekhawat
14 years ago

E1 = ALL 4 R WHITE(WWWW)

E2= 3 WHITE 1 BLACK            

E3 = 2 W AND 2 Black

each of these events hve equal prob.=1/3

let the event of withdrawing 2 white balls be A

now P(A/E1)=2C2/4C2

P(A/E2)=3C2/4C2

P(A/E3)=4C2/4C2

now apply bayes theorem

P(E3/A)= (P(E3)*P(A/E3))    /    P(E3)P(A/E3) + P(E2)P(A/E2)  + P(E1)P(A/E1)

THE ANS IS 3/5